# Iterative execution of discrete and inverse discrete Fourier transforms with applications for signal denoising via sparsification

### v3

@@ Abstract
${}^{1}$epartment of Biomedical Data Science
@@ 1 Problem statement
$\mathbf{x}_{k}=\begin{cases}h(\mathbf{x}_{k})\ &k=0\\ h(dft^{-1}(g(dft(\mathbf{x}_{k-1}))))&k\geq 1\\ \end{cases}$ (3)
@@
$i=1$triangleright Initialize iteration index
2:$\mathbf{x}_{0}=\mathbf{x}$triangleright Initialize $\mathbf{x}_{i}$
@@
4:     $\mathbf{x}^{*}_{i}=h(\mathbf{x}_{i-1})$triangleright Apply function $h():\mathbb{R}^{n}\to\mathbb{R}^{n}$ to $\mathbf{x}_{i-1}$
@@
[…]
8:     $\mathbf{w}^{*}_{i}=g(\mathbf{w}_{i})$triangleright Apply function $g():\mathbb{C}^{n}\to\mathbb{C}^{n}$ to complex vector $\mathbf{w}_{i}$
[…]
10:     $i=i+1$triangleright Increment iteration index return $(\mathbf{x}^{*}_{i},i)$triangleright Return the output of the final execution of $h()$
@@ 1.2 Generalizations
[…]
@@ 2 Related techniques
Algorithm (1) is related to both standard discrete Fourier analysis techniques and iterative algorithms that alternate between dual problem representations. Although a detailed comparison is not possible without defining $h()$, $g()$, and $c()$, some general observations are possible.
@@ 4 Detection of spike signals using iterative convergence
  • Define $g()$ to generate a sparse $\mathbf{w}$ where all elements $|w_{i}|\leq 1/n\sum_{j=1}^{n}|w_{j}|$ are set to $0+0i$ (here the ||||| | operation represents the magnitude of the complex number $w_{j}$).
@@ 4.2 Detection of spike signal in matrix-valued input
Figure 7: Output of the IterativeFT method on a matrix simulated according to Section 4.2. The top panels show the $\mathbf{S}$ and $\mathcal{E}$ matrices. The remaining panels show the input and output of $h()$ after each iteration with the error relative to $\mathbf{S}$ quantified as MSE.
To explore the generalization where the input is a matrix rather than a vector, we also tested a simulation model where an input $n\times n$ matrix $\mathbf{X}$ is generated as follows:
@@ 5.1 Simulation design
    Alternating sign spike: The sparsity pattern is similar to the uniform spike model above but the non-zero values alternate between $\delta$ and $-\delta$. The first panel in Figure 11 below illustrates this signal pattern.
@@
For the simulation results presented in Sections 5.3-5.5 below, $n=1024$ and $\delta=2$. The noise variance, $\sigma^{2}$, was set to yield a signal-to-noise ratio ranging between 0.05 and 2. The $n$ value of 1024 corresponds to a Nyquist frequency of 512 Hz and the spike period $\lambda$ was set to yield a ratio of spike frequency to Nyquist frequency ranging between 0.0625 (for $\lambda=32$) and 1 (for $\lambda=2$). For the single examples in Figures 8, 11 and 14, $n=128$, $\lambda=8$ (i.e., a ratio of spike frequency to Nyquist frequency of 0.25) and $\sigma^{2}$ was set to yield a signal-to-noise ratio of 1.
@@ 5.2 Comparison methods
  • Butterworth bandpass filtering: This method is realized by a Butterworth [11] 3-order bandpass filter (as implemented by the butter() function in v0.3-5 of the gsignal R package [12]) where the band is set to the spike signal frequency (as a fraction of the Nyquist frequency) ±$\pm 0.05$, e.g, if the spike signal is 0.25 of the Nyquist frequency, the band is set to 0.2 to 0.3.
@@ 5.3 Results for uniform spike model
Figures 8-10 illustrate the comparative denoising performance for the evaluated techniques on data simulated according to the uniform spike model. Figure 8 shows the outputs for a single example vector with a signal-to-noise (SNR) ratio of 1 and ratio of spike signal frequency to Nyquist frequency of 0.25. For this example, the IterativeFT method is able to perfectly recover the spike signal. In contrast, the other evaluated techniques have only marginal signal recovery performance.
@@
Figure 10 illustrates the relationship between denoising performance and the SNR value at a spike frequency that is 0.25 of the Nyquist frequency. Similar to Figure 9, both of the simple thresholding methods have performance that is only marginally better than null but, as expected, does improve slightly increasing SNR. The IterativeFT method is substantially better than all comparison techniques at SNR values above 0.5. At low SNR values, performance of the IterativeFT method decreases and falls below both the wavelet and Butterworth filtering techniques at SNR values below $\sim$0.25. The relative performance of the wavelet and Butterworth methods shows a steady decline with increasing SNR, which is surprising the error should decrease as SNR increases. This unexpected trend is due to the fact that the relative rather than absolute MSE is shown. In particular, the absolute MSE for these methods does decrease as the SNR increases, however, the MSE relative to no modifications increases with both filtering methods giving worse than null performance at SNR values greater than $\sim$ 1.1.
@@ 5.4 Results for alternating sign spike model
Figures 11-13 illustrate comparative denoising performance on data simulated according to the alternating sign spike model. Figure 11 shows the outputs for a single example vector with a signal-to-noise (SNR) ratio of 1 and ratio of spike signal frequency to Nyquist frequency of 0.25. Similar to the uniform spike example in Figure 8, the IterativeFT method perfectly recovers the spike signal with poor performance by the comparative techniques.
@@
Figures 12 and 13 visualize the performance of the evaluated methods on the alternating sign model relative to signal frequency and SNR. The results are generally similar to those for the uniform spike model with the exception that the Butterworth bandpass filter has worse than null performance across all tested signal frequencies.
@@ 5.5 Results for alternating size spike model
Figures 14-16 illustrate comparative denoising performance on data simulated according to the alternating size spike model. Figure 14 shows the outputs for a single example vector with a signal-to-noise (SNR) ratio of 1 and ratio of spike signal frequency to Nyquist frequency of 0.25. Similar to the uniform spike and alternating sign examples, the IterativeFT method provides excellent signal recovery though it does underestimate the magnitude of the smaller spike; the other evaluated methods are worse by comparison.
@@
Figures 15 and 16 visualize the performance of the evaluated methods on the alternating size model relative to signal frequency and SNR. The results are generally similar to those for the uniform spike model with the exception that performance for the Butterworth bandpass filter is never better than the IterativeFT method.

### v4

@@ Abstract
1Department of Biomedical Data Science
@@ 1 Problem statement
[…]
@@
  • $\mathbf{x}\in\mathbb{R}^{n}$ $\triangleright$ Input data
  • $i_{m}$ $\triangleright$ The maximum number of iterations
@@
  • $\mathbf{y}\in\mathbb{R}^{n}$ $\triangleright$ Output data
  • $i_{c}$ $\triangleright$ Number of iterations completed
1:$i=1$ $\triangleright$ Initialize iteration index
[…]
@@
4:  $\mathbf{x}^{*}_{i}=h(\mathbf{x}_{i-1})$ $\triangleright$ Apply function $h():\mathbb{R}^{n}\to\mathbb{R}^{n}$ to $\mathbf{x}_{i-1}$
@@
7:  $\mathbf{w}_{i}=\textit{dft}(\mathbf{x}^{*}_{i})$ $\triangleright$ Compute discrete Fourier transform of $\mathbf{x}^{*}_{i}$
8:  $\mathbf{w}^{*}_{i}=g(\mathbf{w}_{i})$ $\triangleright$ Apply function $g():\mathbb{C}^{n}\to\mathbb{C}^{n}$ to complex vector $\mathbf{w}_{i}$
9:  $\mathbf{x}_{i}=\textit{dft}^{-1}(\mathbf{w}^{*}_{i})$ $\triangleright$ Compute inverse discrete Fourier transform of $\mathbf{w}^{*}_{i}$
10:  $i=i+1$ $\triangleright$ Increment iteration index return $(\mathbf{x}^{*}_{i},i)$ $\triangleright$ Return the output of the final execution of $h()$
@@ 1.2 Generalizations
    Matrix-valued input: Instead of accepting a vector $\mathbf{x}\in\mathbb{R}^{n}$ as input, the algorithm could accept a matrix $\mathbf{X}\in\mathbb{R}^{n\times m}$ (or higher-dimensional array) with $\textit{dft}()$ and $\textit{dft}^{-1}()$ replaced by two-dimensional counterparts.
@@ 2 Related techniques
Algorithm (1) is related to both standard discrete Fourier analysis techniques, the iterative Gerchberg-Saxton (GS) algorithm [5], and iterative algorithms that alternate between dual problem representations. Although a detailed comparison is not possible without defining $h()$, $g()$, and $c()$, some general observations are possible.
@@ 2.1 Relationship to standard discrete Fourier analysis
2.2 Relationship to Gerchberg-Saxton (GS) algorithm
The Gerchberg-Saxton (GS) algorithm [5] is an interative Fourier-based technique that was developed to estimate phase information for electron microscopy data. In the case of electron microscopy data, only the amplitudes of the image and Fourier transform of the image (i.e., the diffraction plane) are measured and the phase information must be estimated. The GS method performs this estimation by initializing the phase information to random values and then iteratively applying forward and inverse discrete Fourier transforms on the image or diffraction plane amplitude data normalized using the most recent estimated phase information and then updating the phase estimates using the known amplitudes. The method terminates when the error between the reconstructed and measured amplitudes falls below a specific threshold.
While the GS method iteratively applies forward and inverse discrete Fourier transforms, it is functionally distinct from the class of methods considered in this paper. Specifically, we are considering methods that only accept just a single input (e.g., real domain data) and iteratively apply specified functions to the data and Fourier transformation of the data. In contrast, the GS method accepts two inputs that represent the phaseless amplitudes of a dataset and the amplitudes of the Fourier transform of that data and then use the iteration to estimate the missing phase information. To the authors knowledge, the GS algorithm is the only existing iterative Fourier technique prior to this work.
@@ 4 Detection of spike signals using iterative convergence
  • Define $g()$ to generate a sparse $\mathbf{w}$ where all elements $|w_{i}|\leq 1/n\sum_{j=1}^{n}|w_{j}|$ are set to $0+0i$ (here the $||$ operation represents the magnitude of the complex number $w_{j}$).
@@ 4.2 Detection of spike signal in matrix-valued input
Figure 7: Output of the IterativeFT method on a matrix simulated according to the noisy grid design in Section 4.2. The top panels show the $\mathbf{S}$ and $\mathcal{E}$ matrices. The remaining panels show the input and output of $h()$ after each iteration with the error relative to $\mathbf{S}$ quantified as MSE.
To explore the generalization where the input is a matrix rather than a vector, we also tested several 2D simulation models. The first model is a noisy $n\times n$ grid $\mathbf{X}$ generated as follows:
@@
For the second example, we simulated a rank 1 matrix $\mathbf{X}$ with a sinusoidal data pattern as follows:
  • Create a length $100$ vector $t$ as 100 uniformly spaced points between 0 and 1.
  • Generate a length $100$ vector $\mathbf{x}$ that contains the sinusoidal signal $sin(3\pi t^{2})^{2}$.
  • Generate a length $100$ vector $\mathbf{y}$ that contains the sinusoidal signal $sin(\pi t)^{2}$.
  • Create the signal matrix $\mathbf{S}$ as $\mathbf{S}=\mathbf{x}\mathbf{y}^{T}$.
  • Generate an $100\times 100$ noise matrix $\mathcal{E}$ whose elements are independent $\mathcal{N}(0,0.1)$ random variables.
  • Create the input $100\times 100$ matrix $\mathbf{X}$ as $\mathbf{X}=\mathbf{S}+\mathcal{E}$.
Figure 8: Output of the IterativeFT method on a matrix simulated according to rank 1 sinusoidal model in Section 4.2. The top panels show the $\mathbf{S}$, $\mathcal{E}$, and $\mathbf{X}$ matrices. The remaining panels show the input and output of $h()$ after different iterations with the error relative to $\mathbf{S}$ quantified as MSE.
Figure 8 shows an example of $\mathbf{X}$ generated according to this sinusoidal simulation model. For this specific example, the IterativeFT method converges in 72 iterations and recovers a sparse version of the sinusoidal input pattern.
@@ 4.3 Analysis of spike signal detection
To understand the simulation-based results displayed in Figures 1, 2, 3, 4, 5, and 6, we can evaluate the mathematical representation of one iteration of the algorithm given by (3) for $k\geq 1$ when $h()$ and $g()$ induce sparsity. To restate (3):
@@ 5.1 Simulation design
    Alternating sign spike: The sparsity pattern is similar to the uniform spike model above but the non-zero values alternate between $\delta$ and $\delta$. The first panel in Figure 12 below illustrates this signal pattern.
@@
For the simulation results presented in Sections 5.3-5.5 below, $n=1024$ and $\delta=2$. The noise variance, $\sigma^{2}$, was set to yield a signal-to-noise ratio ranging between 0.05 and 2. The $n$ value of 1024 corresponds to a Nyquist frequency of 512 Hz and the spike period $\lambda$ was set to yield a ratio of spike frequency to Nyquist frequency ranging between 0.0625 (for $\lambda=32$) and 1 (for $\lambda=2$). For the single examples in Figures 9, 12 and 15, $n=128$, $\lambda=8$ (i.e., a ratio of spike frequency to Nyquist frequency of 0.25) and $\sigma^{2}$ was set to yield a signal-to-noise ratio of 1.
@@ 5.2 Comparison methods
  • Butterworth bandpass filtering: This method is realized by a Butterworth [12] 3-order bandpass filter (as implemented by the butter() function in v0.3-5 of the gsignal R package [13]) where the band is set to the spike signal frequency (as a fraction of the Nyquist frequency) $\pm 0.05$, e.g, if the spike signal is 0.25 of the Nyquist frequency, the band is set to 0.2 to 0.3.
@@ 5.3 Results for uniform spike model
Figures 9-11 illustrate the comparative denoising performance for the evaluated techniques on data simulated according to the uniform spike model. Figure 9 shows the outputs for a single example vector with a signal-to-noise (SNR) ratio of 1 and ratio of spike signal frequency to Nyquist frequency of 0.25. For this example, the IterativeFT method is able to perfectly recover the spike signal. In contrast, the other evaluated techniques have only marginal signal recovery performance.
@@
Figure 11 illustrates the relationship between denoising performance and the SNR value at a spike frequency that is 0.25 of the Nyquist frequency. Similar to Figure 10, both of the simple thresholding methods have performance that is only marginally better than null but, as expected, does improve slightly increasing SNR. The IterativeFT method is substantially better than all comparison techniques at SNR values above 0.5. At low SNR values, performance of the IterativeFT method decreases and falls below both the wavelet and Butterworth filtering techniques at SNR values below $\sim$0.25. The relative performance of the wavelet and Butterworth methods shows a steady decline with increasing SNR, which is surprising the error should decrease as SNR increases. This unexpected trend is due to the fact that the relative rather than absolute MSE is shown. In particular, the absolute MSE for these methods does decrease as the SNR increases, however, the MSE relative to no modifications increases with both filtering methods giving worse than null performance at SNR values greater than $\sim$ $1.$
@@ 5.4 Results for alternating sign spike model
Figures 12-14 illustrate comparative denoising performance on data simulated according to the alternating sign spike model. Figure 12 shows the outputs for a single example vector with a signal-to-noise (SNR) ratio of 1 and ratio of spike signal frequency to Nyquist frequency of 0.25. Similar to the uniform spike example in Figure 9, the IterativeFT method perfectly recovers the spike signal with poor performance by the comparative techniques.
@@
Figures 13 and 14 visualize the performance of the evaluated methods on the alternating sign model relative to signal frequency and SNR. The results are generally similar to those for the uniform spike model with the exception that the Butterworth bandpass filter has worse than null performance across all tested signal frequencies.
@@ 5.5 Results for alternating size spike model
Figures 15-17 illustrate comparative denoising performance on data simulated according to the alternating size spike model. Figure 15 shows the outputs for a single example vector with a signal-to-noise (SNR) ratio of 1 and ratio of spike signal frequency to Nyquist frequency of 0.25. Similar to the uniform spike and alternating sign examples, the IterativeFT method provides excellent signal recovery though it does underestimate the magnitude of the smaller spike; the other evaluated methods are worse by comparison.
@@
Figures 16 and 17 visualize the performance of the evaluated methods on the alternating size model relative to signal frequency and SNR. The results are generally similar to those for the uniform spike model with the exception that performance for the Butterworth bandpass filter is never better than the IterativeFT method.